tire string is 2.00 m long a mass of 400 g. the vibrating section of the string is 1.90 m long. a. What tension is needed to tune this string properly b. What are the next two harmonis that can play on this tring.
Answer #1

(a) Mass density µ = 0.400/2.00 = 0.200kg/m

Now f = v/2L = sqrt(T/µ)/2L

So T = 4(L*f)^2*µ = 4*(1.90*27.5)^2*0.200 = 2184 N

(b) Given, f1 = 27.5 Hz

Next two harmonics are:

f2 = 2f1 = 55 Hz

f3 = 3f1 = 82.5 Hz